OpenStudy (anonymous):

Find the equation of the circle whose center and radius are given. center ( 5, 6), radius = 3

OpenStudy (anonymous):

(x-5)^2+(y-6)^2=9

OpenStudy (anonymous):

Johny, the equation of a circle is this: (x-h)^2+(y-k)^2=r^2 where h is the x component of the center of the circle and k is the y component. r is the radius of course.